OpenStudy (anonymous):

Alfred is practicing typing. The first time he tested himself, he could type 26 words per minute. After practicing for a week, he could type 28 words per minute. After two weeks he could type 30 words per minute. Based on this pattern, predict how fast Alfred will be able to type after four weeks of practice.

OpenStudy (anonymous):

34

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

it is 30+2+2= 30+4= 34